If the area of the rectangle is x2 +7x +12 square meters and the length and width of a second rectangle are 3 meters longer, what is the area of the second rectangle?
Answer by stanbon(75887)

If the area of the rectangle is x2 +7x +12 square meters and the length and width of a second rectangle are 3 meters longer, what is the area of the second rectangle?
----------
Original area = x^2 +7x + 12 = (x+3)(x+4)
-----
New area = (x+3+3)(x+4+3) = (x+6)(x+7) = x^2+13x+42
==========================
Cheers,
Stan H.